if a rectangle has a diagonal of 17 and a width of 5 what is the area?
Answer by drk(1908)

We need Pythagorean theorem as
a^2 + b^2 = c^2 and we get
5^2 + b^2 = 17^2
25 + b^2 = 289
b^2 = 264
take a square root to get
b =16
---
a = base x height
a = 16 x 5 = 80